If you had an investment opportunity that promises to pay you $29,000 in four years and you could earn a 8% annual return investing your money elsewhere What is the most you should be willing to invest today in this opportunity? (FV of $1, PV of $1, FVA of $1, and PVA of $1). (Use appropriate factor(s) from the tables provided. Round final answer to the nearest whole dollar.)
Answer #1
Maximum investment = 29000/FV of $1(i=8%n=4) = 29000/1.360489= $21316